Need to buy $65 worth of a liquid. Can only buy 40 bottles. One bottle costs $1.80 and another costs $1.25. Need to buy the most $1.80 as possible. Need an equation. Thx.
Answer by stanbon(75887)

Need to buy $65 worth of a liquid. Can only buy 40 bottles. One bottle costs $1.80 and another costs $1.25. Need to buy the most $1.80 as possible. Need an equation.
---------------------
Let x be # of 1.80 bottles: Let y be # of 1.25 bottles.
---------------------------
Quantity Equation: x+y = 40
Value Equation: 1.8x + 1.25y <= 65
--------------------------------------
x = 40-y
1.8(40-y) + 1.25y <= 65
72 - 0.55y = 65
0.55y = 7
y = 12.72 (number of 1.25 bottles: rounded down to 12 bottles
The x = 28 when rounded up (# of $1.80 bottles)
=====================
Cheers,
Stan H.